# Exploring Composition Mixing in Kilonova Ejecta with Ray-by-ray Simulations

### v1

@@ I Introduction
Numerical relativity (NR) simulations play an essential role in understanding the ejecta of BNSMs. They robustly predict that BNSM merger ejecta are composed of two components: dynamical and disk ejecta. The dynamical ejecta are generated on dynamical timescales. They consist of a component ejected by tidal torques [Rosswog1999, Radice2016, Dietrich2017, Bernuzzi2020, Nedora2021], and a component launched by shocks during and after the merger [Hotokezaka2013, Bauswein2013, Sekiguchi2016, Dietrich2017, Radice2018]. The disk ejecta is produced by the winds from mergers on a longer timescale [Siegel2017, Siegel2018, Fujibayashi2018, Fernandez2019, Shibata2019, Nedora2019]. The tidal ejecta are cold and neutron-rich, with electron fractions $Y_{e}\sim 0.1$ ($Y_{e}$ is the ratio between the net number of electrons and positrons, and baryons) [Rosswog1999, Goriely2011]. The shock ejecta are heated by shocks and reprocessed by neutrino irradiation to higher $Y_{e}$ [Wanajo2014, Radice2016], especially in the polar regions with higher neutrino flux. A low-$Y_{e}$ ($\lesssim 0.25$) equatorial component, often found in the dynamical ejecta and late disk ejecta, produces lanthanides and actinides by $r$-process, leading to high opacity [Kasen2013]. It powers kilonovae whose light curves peak in the infrared with timescales of days (red kilonovae; [Metzger2014, Perego2017]). In the polar dynamical ejecta and most disk ejecta, higher $Y_{e}$ values suppress the synthesis of lanthanides and actinides. Consequently, this material retains low opacity, powering kilonovae that peak in the UV/optical with timescales of hours (blue kilonovae; [Metzger2014, Perego2014, Martin2015, Metzger2019, Tanaka2017, Curtis2024]).
The evolution of the ejecta during the kilonova phase is challenging to model in the same way as BNSMs, given the orders-of-magnitude difference between the merger and kilonova timescales and lengthscales. Traditional models of kilonova evolution often assume homologously expanding ejecta without radiation-hydrodynamic evolution (e.g., [Li1998, Metzger2010, Lippuner2015, Perego2022]). Some models implement Monte Carlo radiative transfer [Kasen2017, Wollaeger2018, Bulla2019, Korobkin2021]. Nuclear heating rates are typically derived from fits to nucleosynthesis calculations [Wu2022, Ricigliano2024]. Recent works simulate the radiation-hydrodynamics of the ejecta, improving the accuracy of ejecta evolution [Roberts2011, Rosswog2014, Montes2016, Wu2022, Just2023]. The use of nuclear networks (NNs) sheds light on nucleosynthesis yields from BNSMs [Ricigliano2024] and on full NN calculations on the heating rate [Lippuner2015, Lippuner2017]. The coupling of hydrodynamics and in-situ NNs improves the fidelity of the abundances and light curves predictions [Magistrelli2024, Magistrelli:2025xja].
@@ II.1 Initial Profiles
Angular profiles are extracted from NR simulations of an equal-mass BNS merger with a gravitational mass of $1.35M_{\odot}$ per star [Radice2023]. The merger was evolved by the NR THC code [Radice2012, Radice2014a, Radice2014b], with an M1 neutrino transport extension [Radice2022, Zappa2023] and DD2 EoS [Hempel2010, Typel2010]. The simulation ran at a spatial resolution (SR; $\Delta x=185\ \mathrm{m}$) to 100 ms post-merger, and we extract angular profiles at $R_{\mathrm{ext}}=295\ \mathrm{km}$. We collect unbound material according to the Bernoulli criterion [Foucart2021], resulting in a total mass of the ejecta of $7.38\times 10^{-3}M_{\odot}$. Using the geodesic criterion to determine the dynamical ejecta [Nedora2021, Combi2023, Ricigliano2024], we identify a dynamical component of $2.69\times 10^{-3}M_{\odot}$, and a disk wind component of $4.68\times 10^{-3}M_{\odot}$. The dynamical ejecta have a shock-heated component with high entropy, and a tidal component with low entropy. The shock-heated component undergoes more intense neutrino captures, increasing its electron fraction to $Y_{e}>0.1$. Additionally, the dense material in the equatorial plane shields the outer layers of the ejecta from neutrino irradiation, thereby maintaining low electron fractions. In contrast, at higher latitudes, the neutrino flux makes the ejecta more proton-rich [Jacobi2024]. We show the distributions of density, electron fraction, entropy, and temperature in Fig. 1. The outer layers of the ejecta at the equator have low electron fraction and entropy.
We divide the ejecta into 51 azimuthally averaged angular profiles, with the properties of each profile dependent on the azimuthal angle integrated out. The profiles are mapped into the Lagrangian mass coordinates by placing layers on top of the most interior layer $R_{\mathrm{ext}}$. We scale masses of the profiles by a factor of $\lambda_{\theta}=4\pi/\Delta\Omega$, where $\Delta\Omega=2\pi\sin\theta\Delta\theta$ is the solid angle of the profile at the polar angle $\theta$ [Magistrelli2024], with all intensive properties unchanged. The profiles are further rearranged into 500 mass shells, with the finest grids at the inner and outer boundaries, to better resolve sharp density gradients and the rapid expansion. Each profile is then evolved as an independent radiation-hydrodynamics spherically symmetric model. Eventually, we map the 1-D simulations back to the axisymmetric ejecta by multiplying all extensive properties by a factor of $1/\lambda_{\theta}$.
@@ II.2 Radiation-Hydrodynamics
At low temperatures, the ejecta EoS follows the Paczyński EoS [Paczynski1983], which includes ideal ion gas, electron gas with ideal, non-relativistic degenerate and relativistic degenerate components, and ideal photon gas. At high temperatures, we use a tabulated Helmholtz EoS [Timmes2000, Lippuner2017] implemented by [Magistrelli2024]. We use the same analytic opacity as [Wu2022] to calculate the luminosity $L$.
@@ II.4 Composition Mixing Scheme
Rather than explicitly calculating the flux of the current timestep, the implicit Euler method calculates the flux based on the mass fractions of the next timestep. This method maintains stable numerical evolution without requiring a smaller timestep than the hydrodynamic timestep.
@@ III Results
Figure 5: The combined global abundances with solar $r$-process residual abundances from [Prantzos2020], normalized at $A=195$. The axes are defined the same as Fig. 4. The abundances of the mixing and no-mixing models are consistent except for some light elements with $A<50$. The global yield is dominated by the angular rays where specific elements are produced most efficiently, which effectively dilutes the minor abundance differences found in other rays. The abundances of the second- and third-peak elements are comparable to the solar $r$-process abundances, but the first-peak and rare-earth elements are underproduced.
@@ III.2 Abundance Yields
In Fig. 6 and 7, we show the evolution maps of abundances at mass numbers $A=195$ (the third peak) and $A=240$ (actinides), respectively. The $r$-process first produces heavy elements in the outer regions of the ejecta near the equator, where the ejecta have low entropy ($s\lesssim 10k_{\mathrm{B}}/\mathrm{baryon}$, Lippuner2017) at the start of the simulation, and subsequently takes place in neutron-rich mass shells. The spatial distributions of these heavy elements at 1 s, when free neutrons have run out, match the spatial distributions of free neutrons at 0.1 s. As we have seen in the $Y_{e}$ and $Y_{\mathrm{n}}$ distributions in Fig. 2 and 3, mixing influences the rays in the transition between rich and poor free neutron regions, where it smooths the fine-scale structures of the heavy elements, but does not alter the general spatial distribution and total global abundances.
@@ Appendix B Analytic Electron Fraction Profile
Figure 12: Final abundances of the analytic $Y_{e}$ profile. The coordinates are defined the same as Fig. 4. Mixing strongly changes the abundances of elements before the second $r$-process peak, but leaves heavier elements unaffected.
We create an analytic profile for the electron fraction $Y_{e}$ to verify that our mixing scheme works as expected. We artificially redefine the inner $Y_{e}$ to 0.45 and the outer $Y_{e}$ to 0.15, interpolating linearly from the two values between $0.45\sim 0.55M_{\mathrm{total}}$. This prescription provides a sharp transition in this range and improves the resolution of mixing effects. We directly take all other physical properties from the NR profile (see Sec. II.1), and then initialize the NN as described in Sec. II.3.

### v2

@@ I Introduction
Numerical relativity (NR) simulations play an essential role in understanding the ejecta of BNSMs. They robustly predict that BNSM merger ejecta are composed of two components: dynamical and disk ejecta. The dynamical ejecta are generated on dynamical timescales. They consist of a component ejected by tidal torques [Rosswog1999, Rezzolla2010, Radice2016, Dietrich2017, Bovard2017, Bernuzzi2020, Nedora2021], and a component launched by shocks during and after the merger [Hotokezaka2013, Bauswein2013, Sekiguchi2016, Dietrich2017, Bovard2017, Radice2018]. The disk ejecta is produced by the winds from mergers on a longer timescale [Siegel2017, Siegel2018, Fujibayashi2018, Fernandez2019, Shibata2019, Nedora2019, Kiuchi2023]. The tidal ejecta are cold and neutron-rich, with electron fractions $Y_{e}\sim 0.1$ ($Y_{e}$ is the ratio between the net number of electrons and positrons, and baryons) [Rosswog1999, Goriely2011]. The shock ejecta are heated by shocks and reprocessed by neutrino irradiation to higher $Y_{e}$ [Wanajo2014, Radice2016], especially in the polar regions with higher neutrino flux. A low-$Y_{e}$ ($\lesssim 0.25$) equatorial component, often found in the dynamical ejecta and late disk ejecta, produces lanthanides and actinides by $r$-process, leading to high opacity [Kasen2013]. It powers kilonovae whose light curves peak in the infrared with timescales of days (red kilonovae; [Metzger2014, Perego2017]). In the polar dynamical ejecta and most disk ejecta, higher $Y_{e}$ values suppress the synthesis of lanthanides and actinides. Consequently, this material retains low opacity, powering kilonovae that peak in the UV/optical with timescales of hours (blue kilonovae; [Metzger2014, Perego2014, Martin2015, Metzger2019, Tanaka2017, Curtis2024]).
The evolution of the ejecta during the kilonova phase is challenging to model in the same way as BNSMs, given the orders-of-magnitude difference between the merger and kilonova timescales and lengthscales. Traditional models of kilonova evolution often assume homologously expanding ejecta without radiation-hydrodynamic evolution (e.g., [Li1998, Metzger2010, Lippuner2015, Perego2022]). Some models implement Monte Carlo radiative transfer [Kasen2017, Wollaeger2018, Bulla2019, Korobkin2021, Collins2023, Shingles2023]. Nuclear heating rates are typically derived from fits to nucleosynthesis calculations [Wu2022, Rosswog2024, Ricigliano2024]. Recent works simulate the radiation-hydrodynamics of the ejecta, improving the accuracy of ejecta evolution [Roberts2011, Rosswog2014, Montes2016, Wu2022, Just2023, Kawaguchi2023, Kawaguchi2024]. The use of nuclear networks (NNs) sheds light on nucleosynthesis yields from BNSMs [Bovard2017, Ricigliano2024] and provides high-fidelity estimates of the heating rate [Lippuner2015, Lippuner2017]. The coupling of hydrodynamics and in-situ NNs improves the fidelity of the abundances and light curves predictions [Magistrelli2024, Magistrelli:2025xja].
@@ II.1 Initial Profiles
Angular profiles are extracted from NR simulations of an equal-mass BNS merger with a gravitational mass of $1.35M_{\odot}$ per star [Radice2023]. The merger was evolved by the NR THC code [Radice2012, Radice2014a, Radice2014b], with an M1 neutrino transport extension [Radice2022, Zappa2023] and DD2 EoS [Hempel2010, Typel2010]. The simulation ran at a spatial resolution ($\Delta x=185\ \mathrm{m}$) to 100 ms post-merger, and we extract angular profiles at $R_{\mathrm{ext}}=295\ \mathrm{km}$. The resolution of the NR simulation in the region covering the NSs before merger and the remnant is of $1.85\times 10^{4}$ cm. The resolution in the region where we extract ejecta data is of $5.9\times 10^{5}$ cm. We collect unbound material according to the Bernoulli criterion [Foucart2021], resulting in a total mass of the ejecta of $7.38\times 10^{-3}M_{\odot}$. Using the geodesic criterion to determine the dynamical ejecta [Nedora2021, Combi2023, Ricigliano2024], we identify a dynamical component of $2.69\times 10^{-3}M_{\odot}$, and a disk wind component of $4.68\times 10^{-3}M_{\odot}$. The dynamical ejecta have a shock-heated component with high entropy, and a tidal component with low entropy. The shock-heated component undergoes more intense neutrino captures, increasing its electron fraction to $Y_{e}>0.1$. Additionally, the dense material in the equatorial plane shields the outer layers of the ejecta from neutrino irradiation, thereby maintaining low electron fractions. In contrast, at higher latitudes, the neutrino flux makes the ejecta more proton-rich [Jacobi2024]. We show the distributions of density, electron fraction, entropy, and temperature in Fig. 1. The outer layers of the ejecta at the equator have low electron fraction and entropy.
We divide the ejecta into 51 azimuthally averaged angular profiles, with the properties of each profile dependent on the azimuthal angle integrated out. The profiles are mapped into the Lagrangian mass coordinates by placing layers on top of the most interior layer $R_{\mathrm{ext}}$. We scale masses of the profiles by a factor of $\lambda_{\theta}=4\pi/\Delta\Omega$, where $\Delta\Omega=2\pi\sin\theta\Delta\theta$ is the solid angle of the profile at the polar angle $\theta$ [Magistrelli2024], with all intensive properties unchanged. The profiles are further rearranged into 500 mass shells, with the finest grids at the inner and outer boundaries, to better resolve sharp density gradients and the rapid expansion. The initial radial location of the outer boundary varies between rays, ranging from $8.54\times 10^{7}$ cm to $1.37\times 10^{8}$ cm. The finest and coarsest radial grid spacings are $5.08\times 10^{3}$ cm and $2.39\times 10^{6}$ cm, respectively. Each profile is then evolved as an independent radiation-hydrodynamics spherically symmetric model. Eventually, we map the 1-D simulations back to the axisymmetric ejecta by multiplying all extensive properties by a factor of $1/\lambda_{\theta}$.
@@ II.2 Radiation-Hydrodynamics
At low temperatures, the ejecta EoS follows the Paczyński EoS [Paczynski1983], which includes ideal ion gas, electron gas with ideal, non-relativistic degenerate and relativistic degenerate components, and ideal photon gas. At high temperatures, we use a tabulated Helmholtz EoS [Timmes2000, Lippuner2017] implemented by [Magistrelli2024]. We use the same analytic opacity as [Wu2022] to calculate the luminosity $L$. The flux density is calculated as blackbody radiation from the photosphere and effective heating in the cells above the photosphere (see Eq. (14) of [Wu2022]). Then, we convolve the flux density with filter functions to obtain light curves.
@@ II.4 Composition Mixing Scheme
Rather than explicitly calculating the flux of the current timestep, the implicit Euler method calculates the flux based on the mass fractions of the next timestep. This method maintains stable numerical evolution without requiring a smaller timestep than the hydrodynamic timestep. The gradient $\nabla X_{i,j}^{k+1}$ is evaluated at cell faces using the mass fractions at the two adjacent cell centers. The face fluxes are then used to compute the divergence $\nabla\cdot(\lambda_{j}^{k+1}D_{j}^{k+1}\nabla X_{i,j}^{k+1})$ at each cell center.
@@ III Results
Figure 5: The combined global abundances with solar $r$-process residual abundances from [Prantzos2020], normalized at $A=195$. The axes are defined the same as in Fig. 4. The abundances of the mixing and no-mixing models are consistent except for some light elements with $A<50$. The global yield is dominated by the angular rays where specific elements are produced most efficiently, which effectively dilutes the minor abundance differences found in other rays. The abundances of the second- and third-peak elements are comparable to the solar $r$-process abundances, but the first-peak and rare-earth elements are underproduced.
@@ III.2 Abundance Yields
In Fig. 6 and 7, we show the evolution maps of abundances at mass numbers $A=195$ (the third peak) and $A=240$ (actinides), respectively. The $r$-process first produces heavy elements in the outer regions of the ejecta near the equator, where the ejecta have low entropy ($s\lesssim 10k_{\mathrm{B}}/\mathrm{baryon}$, [Lippuner2017]) at the start of the simulation, and subsequently takes place in neutron-rich mass shells. The spatial distributions of these heavy elements at 1 s, when free neutrons have run out, match the spatial distributions of free neutrons at 0.1 s. As we have seen in the $Y_{e}$ and $Y_{\mathrm{n}}$ distributions in Fig. 2 and 3, mixing influences the rays in the transition between rich and poor free neutron regions, where it smooths the fine-scale structures of the heavy elements, but does not alter the general spatial distribution and total global abundances.
@@ Appendix B Analytic Electron Fraction Profile
Figure 12: Final abundances of the analytic $Y_{e}$ profile. The coordinates are defined the same as in Fig. 4. Mixing strongly changes the abundances of elements before the second $r$-process peak, but leaves heavier elements unaffected.
We create an analytic profile for the electron fraction $Y_{e}$ to verify that our mixing scheme works as expected. This profile is an extreme case that is unrealistic in BNSM simulations. We artificially redefine the inner $Y_{e}$ to 0.45 and the outer $Y_{e}$ to 0.15, interpolating linearly between these two values over the mass range $0.45\leq M/M_{\mathrm{total}}\leq 0.55$. This prescription provides a sharp transition in this range and improves the resolution of mixing effects. We directly take all other physical properties from the NR profile (see Sec. II.1), and then initialize the NN as described in Sec. II.3.